OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 2 points) Find the 6th partial sum of summation of 7 open parentheses 4 close parentheses to the I minus 1 power from 1 to infinity 78,432 11,204 9,555 2,387

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}7(4)^i-1\]

OpenStudy (anonymous):

@imer can you please help me again?

OpenStudy (imer):

sure, please post.

OpenStudy (anonymous):

the question above...

OpenStudy (imer):

Same concept, this time its the summation from i=1 to 6

OpenStudy (anonymous):

so, it should be 2,387 right?

OpenStudy (imer):

\[\sum_{i=1}^{6}7(4)^{i-1}=7+28+112+448+1792+7168=9,555\]

OpenStudy (anonymous):

mmmmhhhhh!! okkayy!! (:

OpenStudy (anonymous):

do you think you would be able to help me with like 3 more? @imer

OpenStudy (imer):

sure please post.

OpenStudy (anonymous):

Identify whether the series \[\sum_{i=1}^{\infty}16(5)^i -1\] is a convergent or divergent geometric series and find the sum, if possible. This is a convergent geometric series. The sum cannot be found. This is a divergent geometric series. The sum cannot be found. This is a convergent geometric series. The sum is –4. This is a divergent geometric series. The sum is –4

OpenStudy (anonymous):

again the -1 is next to the small i

OpenStudy (imer):

The answer should be "This is a divergent geometric series. The sum cannot be found."

OpenStudy (anonymous):

A skier has decided that on each trip down a slope, she will do 2 more jumps than before. On her first trip she did 6 jumps. Derive the sigma notation that shows how many total jumps she attempts from her fourth trip down the hill through her twelfth trip. Then solve for how many total jumps she attempts from her second trip down the hill through her tenth trip.

OpenStudy (imer):

"Divergent" is when you don't reach a specific value in the limit, in this case it reaches infinity because the sum is getting larger and larger and it will continue to grow forever.\[\sum_{i=1}^{\infty}16(5)^{i-1}=16+80+400+2000+10000+50000 ...\]

OpenStudy (imer):

The question says that in her 1st trip means i=1 she made 6 jumps and those jumps will increase by addition of 2 in each subsequent trips. Now tell me, how many jumps she made in 2nd trip?

OpenStudy (anonymous):

these are tha answer choices: A. \[\sum_{1=2}^{10}2i+4; 144 jumps\] B. \[\sum_{i=2}^{10}2i+4; 180 jumps\] C. \[\sum_{i=2}^{10}4i+2; 234 jumps\] D.\[\sum_{i=2}^{10}4i+2 ;306 jumps\]

OpenStudy (anonymous):

she made 4 jumps right?

OpenStudy (imer):

in her 1st trip,, she made 6 jump, and in 2nd trip she is doing 2 more jumps than before means 8 jumps.

OpenStudy (anonymous):

oh see thats what i thought but someone told me i was wrong last time!

OpenStudy (anonymous):

i think the answer is 234 jumps....

OpenStudy (imer):

and now you can go with the pattern\[6,8,10,12,14,16,18,20...\]and so on. Now lets make a summation formula, shall we?

OpenStudy (anonymous):

oh crap, im running out of time. Im trying to figure it out....

OpenStudy (imer):

Answer is "A"

OpenStudy (anonymous):

thank you!! (: